OpenStudy (anonymous):

fatorize completely y2-x2-6y+6x

OpenStudy (fibonaccichick666):

alright so this is a complete the square problem. The first thing we want to do is group like terms. \[y^{2}-x^{2}-6y+6x\] becomes \[(y^{2}-6y)+(-x^{2}+6x)\] So now I would like you to tell me a number that adds to -6 but multiplies to zero.

OpenStudy (anonymous):

that would be positive 6

OpenStudy (anonymous):

am i correct?

OpenStudy (fibonaccichick666):

I'm sorry I phrased the question incorrectly. 2 numbers that add to -6 but multiply to 0.

OpenStudy (anonymous):

i don't undertsand how to find those numbers. I do not know how to do a complete the square problem.

OpenStudy (anonymous):

But when i look at the original expression, i see a perfect square method of factorizing as well as a comon factor method of factorizing .

OpenStudy (fibonaccichick666):

ok well let's right this as two expressions to find two numbers that add to -6 we take \[x+y=-6\] and to find two numbers that multiply to zero \[xy=0\]

OpenStudy (fibonaccichick666):

so take the multiplication and tell me what could x or y be equal to?

OpenStudy (anonymous):

x+y=-6 y=(-6-x) and x+y=-6 x=(-6-y)

OpenStudy (fibonaccichick666):

and yes there is a common factor method but learning how to complete the square is important so i would like to do it from that perspective; however, you will arrive at the same conclusion.

OpenStudy (anonymous):

okay then, i am willing to learn that method so you can go ahead and teach it to me

OpenStudy (anonymous):

Were my answers correct?

OpenStudy (fibonaccichick666):

ok, so as before, we are looking for two numbers let's call them a and b. we know two pieces of information about these two variables. 1) they multiply to zero ie: \[ab=0\]

OpenStudy (fibonaccichick666):

yes you solved those correctly you can use them for this next part

OpenStudy (fibonaccichick666):

2) they add to -6. ie: \[a+b=-6\]

OpenStudy (fibonaccichick666):

can you solve the system of equations and tell me the value of a and the value of b?

OpenStudy (anonymous):

okay then a+b =-6 b=(-6-a) and a+b=-6 a=(-6-b)

OpenStudy (fibonaccichick666):

now you have not addressed statement 1. please solve that eq.

OpenStudy (anonymous):

okay then,here goes a=(-6-b) and b=(-6-a), therefore, ab=0 (-6-b)(-6-a)=0 (36 +6a) +(6b+ab)=0 6(6+a) b(6+a)=0 (6+a) (6+b)=0 6+a=0 and 6+b=0 a=-6 b=-6

OpenStudy (anonymous):

am i correct?

OpenStudy (fibonaccichick666):

unfortunately no, you should not need to solve 2. only 1 . let's look at this again from the first eq. if i told you that \[a*b=0\] what could you tell me about a or b

OpenStudy (anonymous):

i can say that one is positve and one is negative

OpenStudy (fibonaccichick666):

can you tell me why? or could you tell me the answer to 8x=0

OpenStudy (anonymous):

the answer to that question is x=0

OpenStudy (fibonaccichick666):

so if you have ab=0 what does either a or b have to be?

OpenStudy (anonymous):

0?

OpenStudy (fibonaccichick666):

yes! one of them has to =0

OpenStudy (anonymous):

pswew!!!!

OpenStudy (anonymous):

so that was statement one completed what about statement two?

OpenStudy (fibonaccichick666):

alright, so pick one to =0 then solve 2 for the other

OpenStudy (anonymous):

if i choose a to be equal to zero, then i would be left with b=-6

OpenStudy (fibonaccichick666):

good!!! so if we look to our original problem for the y's , we needed two numbers that add to -6 but multiply to zero. can you fill in the blanks? \[(y+ __ )*(y+ __ )\]

OpenStudy (anonymous):

0, -6

OpenStudy (fibonaccichick666):

=D good can you do the same for the x's? aslo do you know how to FOIL to check your work?

OpenStudy (anonymous):

no i do not know how to foil to check my work

OpenStudy (fibonaccichick666):

ok i can show you that as well, but it will have to wait until later because i have to go to work

OpenStudy (anonymous):

what is FOIL?

OpenStudy (anonymous):

okay then do you use yahoo messenger?

OpenStudy (fibonaccichick666):

ok real quick, it stands for First Inside Outside Last

OpenStudy (fibonaccichick666):

Ok so FOIL-ing it's not bad at all since you are good with variables, I will use those to explain. if you have \[(a+b)*(c+d)\] if you want to simplify you are going to apply FOIL first take the a and distribute it to the (c+d) \[(ac+ad)+(b)*(c+d)\] now distribute the b to the (c+d) this yields \[ac+ad+bc+bd\] And that is all there is to it. ex. 1) \[(x+3)*(x+4)\] 2)\[x^{2}+4x+(3)(x+4)\] 3)\[x^{2}+4x+3x+12\] now combine like terms 4)\[x^{2}+7x+12\] Notice that the b, d terms add to 7 but multiply to 12. Doing this in reverse is called completing the square. =)

OpenStudy (anonymous):

I clearly understand the concept of FOIL ing, after your explanation. I never heard the term until you made mention of it, but i remember calculatiing binomials that result in trinomial expression via FOIL ing. I appreciate your help very much and i wish you well.My concern now is Functions, i do not really understand it but i am reaching there.

OpenStudy (fibonaccichick666):

np and best of luck!